### v1

@@ 3.3.1 Euclidean distance technique
The first method proposed to disentangle intrinsic and LIV delays is based on the computation of the Euclidean distance between the delay versus energy curves at low and high energies. It quantifies the degree of similarity between the shape of two distinct data sets, namely here the time delays in the X-ray and the VHE gamma-ray domains. The Euclidean distance between two data sets $A[i]_{(1\leqslant i\leqslant n)}$ and $B[i]_{(1\leqslant i\leqslant n)}$ containing $n$ points is simply written as follows:
@@
where the denominator is a normalisation term which is introduced for a better graphical representation and readability. The scalar $d_{E}$ possesses two specific identities:
@@
For this multi-wavelength study, two data sets are generated: the first contains synchrotron delays (with inverse Compton delays set to zero), and the other contains inverse Compton delays (with synchrotron delays set to zero). In other words, $A[i]$ (respectively $B[i]$) corresponds to the values of $\Delta t$ below (respectively above) the dip in the SED. To properly apply the method, we make sure there is the same number of non-zero points in each data set, a necessary condition for the method to work as intended and steer relevant information. Then, both sets of delays are extended between $\log_{10}E=-16$ and $\log_{10}E=2$ ($E$ being expressed in TeV), filling the areas where there is no information with zeros. To maximise the sensitivity, we re-scale all the data points with the first non-zero delay value $\Delta t_{\mathrm{first}}$: $\Delta t\longrightarrow\Delta t-\Delta t_{\mathrm{first}}$. Finally, we also need to shift one of the data sets over energy towards the other one to provoke their overlapping. This allows to find the optimal displacement that will minimise the Euclidean distance $d_{E}$. To estimate this optimal displacement, we introduce a modified version of Equation 13 where $d_{E}$ now becomes a function of the displacement $k$:
@@
with $A$ the displaced data set. For the present study, we compute the distance in logarithmic scale by displacing the synchrotron data set ($\equiv A$) over energy by a quantity $\epsilon=10^{k}$ towards the inverse Compton set ($\equiv B$) such that $E_{\mathrm{new}}=E\times 10^{k}$, with $k$ varying from 0 to 16. Although our data sets are discrete, we treat the distributions of time delays as functions and retrieve values with linear interpolations between consecutive points such that $k$ can take any real value.
@@
The Euclidean distance between the synchrotron and inverse Compton data sets is shown in Fig. 6 for three flares illustrating the different intrinsic delay regimes. One can see that in all cases (increasing, flat and decreasing trends), the Euclidean distance $d_{E}$ reaches a clear minimum $d_{E,\mathrm{min}}=d_{E}(k_{\mathrm{min}})$ for intermediate values of $k$, at $k_{\mathrm{min}}$ typically of the order of 9, corresponding to the best match between the shape of the two sets of delays. For small $k$ (here $k<3$), $d_{E}=1$ because the non-zero values of the two sets of delays do not yet overlap. As $k$ increases, data sets non-zero values start to overlap and the Euclidean distance gradually shifts away from 1, towards its minimum value. Once $k$ increases above $k_{\mathrm{min}}$, the similarity between the two data sets and their overlapping area start to decrease again and the Euclidean distance slowly tends towards $d_{E}=1$. One can note also that the specific case of ”flat trend” (in black), which presents little variation of the delays with the frequency, tends to reduce the sensitivity on the Euclidean distance method, reaching a smooth minimum of $d_{E,\mathrm{min}}\sim 0.42$ at $k\sim 9.2\equiv k_{\mathrm{min}}$. Euclidean distance functions for the decreasing and increasing trends reach significantly lower values ($d_{E,\mathrm{min}}\sim 0.25$ and $d_{E,\mathrm{min}}\sim 0.20$ respectively) than for the flat trend case. This is due to the amplified variation of time delays with the energy, leading to a minimum in the Euclidean distance function which is better defined.
All flares with intrinsic time delays significantly increasing or decreasing at high energies show a small and well defined minimum Euclidean distance $d_{E,\mathrm{min}}$, however without reaching a perfect match ($d_{E,\mathrm{min}}=0$) between their synchrotron and inverse-Compton delays. Simulations of several flares covering a large space of parameters extended around the reference set of Table 1 show that the optimal displacement $k$ is always such that $108<k_{\mathrm{min}}<$ 10 and that the minimum Euclidean distance is generally $d_{E,\mathrm{min}}<0.5$ and always such that $d_{E,\mathrm{min}}<0.6$. The maximal value of 0.6 is reached for some flares with time delays following a flat trend.
Figure 6: Euclidean distance functions computed between intrinsic time delays from the synchrotron and inverse Compton domains for the three cases presented in Fig. 2. Each case yields a clear minimum defining the optimal displacement $k_{\mathrm{min}}$ for which we obtain the best match between the two sets of delays. Blue: increasing trend. Black: flat trend. Red: decreasing trend.
@@
Since LIV mostly affects the inverse Compton domain at very high energies while leaving the synchrotron domain unaltered, its contributions tend to decorrelate synchrotron and inverse Compton sets of time delays. As a consequence, the presence of LIV delays can significantly modify the initial pattern of the Euclidean distance function. As illustrated by several cases shown in Fig. 7, the distance function minimum $d_{E,\mathrm{min}}$ can drastically increase when $\tau$ gets different from 0, and the optimal displacement $k_{\mathrm{min}}$ can be modified as a consequence of highly decorrelated sets of data. There are however some specific cases where the introduction of LIV leaves the minimum Euclidean distance almost unchanged as illustrated by a few cases in Fig. 7 (for example, bottom row, right panel, for $\tau=+200$ s TeV-1). Nevertheless, as the introduction of LIV tends to decorrelate the time delays, it always yields either larger or equal values of $d_{E,\mathrm{min}}$.
@@
Assuming next generation of instruments will be able to observe bright flares with a good temporal and large spectral coverage and to detect non-zero time delays, then the Euclidean distance function will become an observable. Fitting it with a hybrid ”leptonic + LIV” model should make possible, in favourable cases, to discriminate between the two origins of the delays. However this will require very high quality sets of data. As a first step, simpler criteria could be applied, for instance taking advantage of the constraints we found above on $d_{E,\mathrm{min}}$ and $k_{\mathrm{min}}$ (for $\tau=0)$. The detection of some $d_{E,\mathrm{min}}>0.6$, $k_{\mathrm{min}}<8$, or $k_{\mathrm{min}}>10$ coming from an intrinsic SSC flare would emphasise the presence of additional non-intrinsic delays possibly due to LIV.
@@ 3.3.2 Hardness-intensity diagrams and hysteresis patterns
The degree of similarity between synchrotron and inverse Compton domains can also be quantified directly from SEDs. Hardness-intensity diagrams (HIDs), showing the time evolution of the SED index (hardness) as a function of the SED flux (intensity), can give rise to specific hysteresis patterns. Such patterns have been found in observed data as presented for instance in Tramacere et al. (2009); Abeysekara et al. (2017); Takahashi et al. (1996), but also simulated with AGN models for instance in Böttcher & Baring (2019). They should ultimately help to constrain emission scenarios. The study of hysteresis patterns appears to be complementary to the Euclidean distance technique and can help consolidate this approach for discrimination between intrinsic and LIV-induced effects in time delays.
@@
To produce HIDs, the hardness is computed over two small energy windows which need to be the same for all SEDs, one in the synchrotron bump and the other in the inverse Compton bump. For the purpose of the present study, the inverse Compton window should be at energies as high as possible. The intensity plotted in the diagram is taken as the mean flux within the energy window, which should be relatively small for the intensity to vary as little as possible. Finally, the window should not overlap with bumps’ maxima to prevent a change of sign of the spectral index. We therefore define the inverse Compton (resp. synchrotron) window as follows: the lower boundary is set to the position of the inverse Compton (resp. synchrotron) peak at the highest energy reached during the flare, and the width is set to one decade in energy.
Figure 8: Hardness-intensity diagrams (HIDs) highlighting hysteresis patterns for X-ray (blue) and gamma-ray (magenta) domains for intrinsic effects only, namely for the three cases with $\tau=0$ shown in Fig. 7. Left: increasing trend, counter-clockwise hysteresis loop at high fluxes. Centre: flat trend, almost no hysteresis. Right: decreasing trend, clockwise hysteresis loop at high fluxes.
The hysteresis pattern for the three flares of Fig. 7 illustrating the different trends of intrinsic delays can be found in Fig. 8. The X-ray domain hysteresis loop is much wider than the gamma-ray one, as expected, since the flux amplitude variation in the SED is more important in the synchrotron domain than in the inverse Compton one. However, the X-ray and gamma-ray hysteresis always follow the same loop orientation at high fluxes. The loop orientation depends on the nature of the flare, and changes at high fluxes according to its intrinsic delay regime. Flares with decreasing (resp. increasing) intrinsic trend show a clockwise (resp. counter-clockwise) hysteresis loop at high X-ray and gamma-ray fluxes. Flares with a flat intrinsic trend show almost no hysteresis, the rising and decaying parts of the flare following the same paths on the HID at high fluxes, as can be expected when there is no intrinsic time delays. Indeed, observing an hysteresis pattern at high flux in a data set means there also should be non-zero intrinsic time delays in the SSC scenario, therefore information on the X-ray domain hysteresis can predict the existence or the absence of intrinsic time delays in the VHE gamma-ray domain in the frame of standard physics.
@@
On the contrary to time delays, LIV contribution to the spectra is barely noticeable and could not be resolved directly from observed SEDs. However, as hysteresis patterns are very sensitive to any variation in SEDs, modification due to LIV can be highlighted through HIDs. Figure 9 illustrates this statement by showing the change of the gamma-ray hysteresis pattern for the flat intrinsic trend when additional LIV effects are injected with various values of $\tau$. At low fluxes, the hardness is greatly affected leading to dramatically different values, and slowly converges as the flux increases where hysteresis patterns then almost overlap. The same qualitative results hold for the two other increasing and decreasing trends (not shown here). Figure 10 shows the HID for the same flare in the X-ray and gamma-ray ranges, with additional LIV effects of $\tau=+400$ s TeV-1 and $\tau=-400$ s TeV-1, and offers a better view of the resolved pattern at high fluxes. The hysteresis pattern for X-ray energies (blue) is unaffected by LIV as expected and can be used as reference of a flat intrinsic trend at high flux. Conversely, the gamma-ray hysteresis pattern can be strongly modified (here in particular for the superluminal case with $\tau<0$) and in both cases shows at high fluxes a clear change of trend in the total time delays due to LIV effects, with the following behaviour:
  • subluminal LIV effects ($\tau>0$) induce a counter-clockwise orientation of the hysteresis loop, and increasing time delays;
@@
Figure 9: HIDs in the gamma-ray energy domain for the flat trend case of Fig. 8 (centre), with various additional LIV contributions (in colour). In the inset plot, the curves are shifted, emphasising the change in curvature orientation at high flux values.
Figure 10: HID highlighting hysteresis patterns for X-ray (blue) and gamma-ray (magenta) domains for the flat trend case of Fig. 8 (center) with additional LIV effects, namely $\tau=+400$ s TeV-1 (top pannel) and $\tau=-400$ s TeV-1 (bottom pannel), showing that LIV effects can significantly modify the hysteresis pattern and even cause a change in loop orientation in the gamma-ray domain. While the X-ray hysteresis pattern is left unchanged relatively to the case at $\tau=0$ of Fig. 8, a rotation pattern appears in the gamma-ray one (either clockwise or counter-clockwise depending on the sign of $\tau$).
@@ 4 Conclusion and perspectives
This article is centred on energy-dependent time delays, arising as one of the most explored signatures of Lorentz invariance violation in astrophysics in the electromagnetic sector. Intrinsic time delays generated in situ by radiative processes of cosmic sources can interfere with these studies and need to be properly disentangled from propagation effects. It would indeed be necessary to accurately discriminate and quantify each contribution to any detected time delay to provide a legitimate constraint on the energy scale $E_{QG}$, and thus a proper interpretation on the tested quantum gravity models.
@@
We investigate this issue with a time dependent modelling of blazar flares induced by particle acceleration and relying on standard leptonic one-zone scenarios. Such models naturally generate intrinsic energy-dependent time delays whether considering SSC models, with or without adiabatic expansion or compression effects, or additional external inverse-Compton emission. The core of the present work focuses on a multi-wavelength study which highlights a strong correlation and symmetry between intrinsic time delays in the X-ray (unaffected by LIV) and gamma-ray (where LIV effects could arise) energy domains. This correlation is due to the fact that in leptonic scenarios the same lepton population is at the origin of the two radiative processes. Two complementary approaches are then proposed to explore and quantify that correlation especially for the pure SSC case where it is the strongest, based on two observables, namely $(i)$ the Euclidean distance between the energy-dependent time delays at X-ray and very high energies, and $(ii)$ the hardness-intensity diagrams built from spectral energy distributions which show hysteresis patterns in the two domains. From these two methods, the presence of any independent external phenomenon modifying time delays such as LIV should be revealed by the weakening or even breaking of the expected intrinsic X-gamma correlation. Moreover, LIV should induce a systematic effect always affecting the correlation in the same direction, with an amplitude varying as a function of the redshift of the cosmic sources.
@@
The present study illustrates how all basic kinds of one-zone leptonic flares (SSC, with or without EIC and adiabatic effects) induced by particle acceleration can grow in two main regimes at high energies, either ”acceleration-driven” or ”cooling-driven” regimes, depending on the high energy lepton population evolution. The evolution of intrinsic time delays as a function of the energy follow three specific trends, described here as increasing, flat and decreasing, which can be explained from the competition between acceleration and radiative processes. The flat trend corresponds to the thin transition between the two main regimes and is obtained only with a fine-tuning of the model parameters. EIC and adiabatic processes do not modify this general picture. However, their presence influences the nature of the flare compared to a pure SSC scenario, and can even change its regime. Additional EIC radiation increases radiative losses and always pushes the flares towards the decreasing trend and ”cooling-driven” regime at VHE. Conversely, adiabatic expansion slows down particle acceleration processes and pushes the flares towards the increasing trend and the acceleration-driven regime (and the opposite for adiabatic compression). For these reasons and for sake of simplicity we mostly applied here the Euclidean distance and HID techniques to pure SSC flares. LIV effects acting only at VHE can reduce and even delete in some cases the strong correlation expected between the X-ray and VHE gamma-ray emission in leptonic scenarios. For a large domain of flare parameters, a minimum normalised Euclidean distance between delays in X-rays and in gamma-rays ($d_{E,\mathrm{min}}<0.6$) was deduced for SSC models of sources at low redshift (z = 0.03). Therefore, obtaining values of $d_{E,\mathrm{min}}>0.6$ from simultaneous X-ray and VHE observations of SSC flares from such sources would sign the presence of additional delays, possibly due to LIV effects. Similarly, the injection of LIV effects can significantly modify the HID hysteresis patterns at VHE energies, and can even inverse the direction of rotation of the gamma-ray hysteresis curve compared to the X-ray one. As a matter of fact, the simple detection of opposite directions of rotation of the hysteresis curve at high fluxes in X-rays and gamma-rays in SSC flares would sign the presence of non-intrinsic time delays possibly due to LIV.
@@
Euclidean distance and HID approaches may be more or less accurate and conclusive to discriminate between the origins of the delays, depending on the observed events and on the instrumental performance of the present or future detectors in term of sensitivity, spectral and temporal resolution. Anyway they both emphasise the decisive importance of simultaneous X-ray and VHE gamma-ray monitoring of variable blazars and of efficient global multi-wavelength alarm networks to catch both rise and decay phases of strong flares. Only under this condition we will succeed in disentangling possible LIV time delays from intrinsic ones with individual sources.
@@
The two techniques proposed here are closely related but complementary and can be adapted depending on the characteristics of available datasets. In particular, a high activity episode sometimes shows multiple flux peaks, such as in the case of the flare of blazar PKS 2155-304 in 2006 (Aharonian et al., 2007). In these conditions, it is reasonable to assume the multiple peaks will have similar intrinsic properties. The $d_{E}$ technique is based on the analysis of X-ray and gamma-ray light curves in several spectral bands. It should be possible to apply it both to a single peak or to multiple peaks, as long as it is possible to measure the lag as a function of energy in the X-ray and VHE domains, with a high enough accuracy. On the other hand, the HID method directly reflects the evolution of the SEDs and relies on the accurate measurement of the flux and spectral index variation in time. This accuracy will probably be achieved only by analysing individual peaks separately, provided photon statistics are high enough. As of now, a first feasibility study for the use of HIDs with CTA is currently ongoing (Rosales De Leon et al., 2024).

### v2

@@ 3.3.1 Euclidean distance technique
The first method proposed to disentangle intrinsic and LIV delays is based on the computation of the Euclidean distance between the delay versus energy curves at low and high energies. It quantifies the degree of similarity between the shape of two distinct datasets, here the time delays in the X-ray and the VHE gamma-ray domains. The Euclidean distance between the two datasets $A[i]_{(1\leqslant i\leqslant n)}$ and $B[i]_{(1\leqslant i\leqslant n)}$ containing $n$ points is simply written as
@@
where the denominator is a normalisation term, which is introduced for better graphical representation and readability. The scalar $d_{E}$ possesses two specific identities:
@@
For this multi-wavelength study, two datasets are generated. The first contains synchrotron delays (with inverse-Compton delays set to zero) and the other contains inverse-Compton delays (with synchrotron delays set to zero). In other words, $A[i]$ (respectively $B[i]$) corresponds to the values of $\Delta t$ below (respectively above) the dip in the SED. To properly apply the method, we make sure there is the same number of non-zero points in each dataset, a necessary condition for the method to work as intended and steer relevant information. Then, the two sets of delays are extended between $\log_{10}E=-16$ and $\log_{10}E=2$ ($E$ being expressed in TeV), filling the areas where there is no information with zeros. To maximise the sensitivity, we re-scale all the data points with the first non-zero delay value $\Delta t_{\mathrm{first}}$: $\Delta t\longrightarrow\Delta t-\Delta t_{\mathrm{first}}$. Finally, we also need to shift one of the datasets over energy towards the other one to provoke their overlapping. This allows us to find the optimal displacement that will minimise the Euclidean distance $d_{E}$. To estimate this optimal displacement, we introduce a modified version of Equation 13, where $d_{E}$ now becomes a function of the displacement $k$,
@@
with $A$ the displaced dataset. For the present study, we computed the distance in logarithmic scale by displacing the synchrotron dataset ($\equiv A$) over energy by a quantity $\epsilon=10^{k}$ towards the inverse Compton set ($\equiv B$) such that $E_{\mathrm{new}}=E\times 10^{k}$, with $k$ varying from 0 to 16. Although our datasets are discrete, we treated the distributions of time delays as functions and retrieved values with linear interpolations between consecutive points such that $k$ can take any real value.
@@
The Euclidean distance between the synchrotron and inverse Compton datasets is shown in Fig. 6 for three flares illustrating the different intrinsic delay regimes. In all cases (increasing, flat, and decreasing trends), the figure shows that the Euclidean distance $d_{E}$ reaches a clear minimum $d_{E,\mathrm{min}}=d_{E}(k_{\mathrm{min}})$ for intermediate values of $k$, at $k_{\mathrm{min}}$ typically of the order of 9, corresponding to the best match between the shape of the two sets of delays. For small $k$ (here $k<3$), $d_{E}=1$ because the non-zero values of the two sets of delays do not yet overlap. As $k$ increases, datasets of non-zero values start to overlap, and the Euclidean distance gradually shifts away from 1 and towards its minimum value. Once $k$ increases above $k_{\mathrm{min}}$, the similarity between the two datasets and their overlapping area start to decrease again and the Euclidean distance slowly tends towards $d_{E}=1$. We also note that the specific case of flat trend (in black), which presents little variation of the delays with the frequency, tends to reduce the sensitivity on the Euclidean distance method, reaching a smooth minimum of $d_{E,\mathrm{min}}\sim 0.42$ at $k\sim 9.2\equiv k_{\mathrm{min}}$. Euclidean distance functions for the decreasing and increasing trends reach significantly lower values ($d_{E,\mathrm{min}}\sim 0.25$ and $d_{E,\mathrm{min}}\sim 0.20$ respectively) than for the flat trend case. This is due to the amplified variation of time delays with the energy, leading to a minimum in the Euclidean distance function, which is better defined.
All flares with intrinsic time delays significantly increasing or decreasing at high energies show a small and well-defined minimum Euclidean distance $d_{E,\mathrm{min}}$, however, without reaching a perfect match ($d_{E,\mathrm{min}}=0$) between their synchrotron and inverse-Compton delays. Simulations of several flares covering a large space of parameters extended around the reference set of Table 1 show that the optimal displacement $k$ is always such that $108<k_{\mathrm{min}}<$ 10 and that the minimum Euclidean distance is generally $d_{E,\mathrm{min}}<0.5$ and always such that $d_{E,\mathrm{min}}<0.6$. The maximum value of 0.6 is reached for some flares with time delays following a flat trend.
Figure 6: Euclidean distance functions computed between intrinsic time delays from the synchrotron and inverse-Compton domains for the three cases presented in Fig. 2. Each case yields a clear minimum defining the optimal displacement $k_{\mathrm{min}}$ for which we obtain the best match between the two sets of delays. Blue: increasing trend. Black: flat trend. Red: decreasing trend.
@@
Since LIV mostly affects the inverse-Compton domain at very high energies, while leaving the synchrotron domain unaltered, its contributions tend to decorrelate synchrotron and inverse-Compton sets of time delays. As a consequence, the presence of LIV delays can significantly modify the initial pattern of the Euclidean distance function. As illustrated by several cases shown in Fig. 7, the distance function minimum $d_{E,\mathrm{min}}$ can drastically increase when $\tau$ becomes different from 0, and the optimal displacement $k_{\mathrm{min}}$ can be modified as a consequence of highly decorrelated sets of data. There are, however, some specific cases where the introduction of LIV leaves the minimum Euclidean distance almost unchanged, as illustrated by a few cases in Fig. 7 (e.g. bottom row, right panel, for $\tau=+200$ s TeV-1). Nevertheless, as the introduction of LIV tends to decorrelate the time delays, it always yields either larger or equal values of $d_{E,\mathrm{min}}$.
@@
Assuming that the next generation of instruments will be able to observe bright flares with a good temporal and large spectral coverage and to detect non-zero time delays, then the Euclidean distance function will become an observable. Fitting it with a hybrid (leptonic + LIV) model should make it possible, in favourable cases, to discriminate between the two origins of the delays. However this will require very high-quality sets of data. As a first step, simpler criteria could be applied, for instance taking advantage of the constraints we found above on $d_{E,\mathrm{min}}$ and $k_{\mathrm{min}}$ (for $\tau=0)$. The detection of some $d_{E,\mathrm{min}}>0.6$, $k_{\mathrm{min}}<8$, or $k_{\mathrm{min}}>10$ coming from an intrinsic SSC flare would emphasise the presence of additional non-intrinsic delays possibly due to LIV.
@@ 3.3.2 Hardness-intensity diagrams and hysteresis patterns
The degree of similarity between synchrotron and inverse-Compton domains can also be quantified directly from SEDs. Hardness-intensity diagrams (HIDs), showing the time evolution of the SED index (hardness) as a function of the SED flux (intensity), can give rise to specific hysteresis patterns. Such patterns have been found in observed data (e.g. in Tramacere et al., 2009; Abeysekara et al., 2017; Takahashi et al., 1996), but also simulated with AGN models (e.g. in Böttcher & Baring, 2019). They should ultimately help to constrain the emission scenarios. The study of hysteresis patterns appears to be complementary to the Euclidean distance technique and can help consolidate this approach for discrimination between intrinsic and LIV-induced effects in time delays.
@@
To produce HIDs, the hardness is computed over two small energy windows that need to be the same for all SEDs, one in the synchrotron bump and the other in the inverse-Compton bump. For the purpose of the present study, the inverse-Compton window should be at energies as high as possible. The intensity plotted in the diagram is taken as the mean flux within the energy window, which should be relatively small for the intensity to vary as little as possible. Finally, the window should not overlap with the bumps’ maxima to prevent a change in sign of the spectral index. We therefore define the inverse-Compton (resp. synchrotron) window as follows: the lower boundary is set to the position of the inverse-Compton (resp. synchrotron) peak at the highest energy reached during the flare, and the width is set to one decade in energy.
Figure 8: Hardness-intensity diagrams (HIDs) highlighting hysteresis patterns for X-ray (blue) and gamma-ray (magenta) domains for intrinsic effects only, namely for the three cases with $\tau=0$ shown in Fig. 7. Left: Increasing trend, anticlockwise hysteresis loop at high fluxes. Centre: Flat trend, almost no hysteresis. Right: Decreasing trend, clockwise hysteresis loop at high fluxes.
The hysteresis pattern for the three flares of Fig. 7 illustrating the different trends of intrinsic delays can be found in Fig. 8. The X-ray domain hysteresis loop is much wider than the gamma-ray one, as expected since the flux amplitude variation in the SED is more important in the synchrotron domain than in the inverse-Compton domain. However, the X-ray and gamma-ray hysteresis always follow the same loop orientation at high fluxes. The loop orientation depends on the nature of the flare, and changes at high fluxes according to its intrinsic delay regime. Flares with decreasing (resp. increasing) intrinsic trend show a clockwise (resp. anticlockwise) hysteresis loop at high X-ray and gamma-ray fluxes. Flares with a flat intrinsic trend show almost no hysteresis, the rising and decaying parts of the flare following the same paths on the HID at high fluxes, as can be expected when there is no intrinsic time delay. Observing a hysteresis pattern at high flux in a dataset means there should also be non-zero intrinsic time delays in the SSC scenario; therefore, information on the X-ray domain hysteresis can predict the existence or the absence of intrinsic time delays in the VHE gamma-ray domain in the frame of standard physics.
@@
Contrary to time delays, the LIV contribution to the spectra is barely noticeable and cannot be resolved directly from observed SEDs. However, as hysteresis patterns are very sensitive to any variation in SEDs, modification due to LIV can be highlighted through HIDs. Figure 9 illustrates this statement by showing the change in the gamma-ray hysteresis pattern for the flat intrinsic trend when additional LIV effects are injected with various values of $\tau$. At low fluxes, the hardness is greatly affected leading to dramatically different values, and slowly converges as the flux increases where hysteresis patterns then almost overlap. The same qualitative results hold for the two other increasing and decreasing trends (not shown here). Figure 10 shows the HID for the same flare in the X-ray and gamma-ray ranges, with additional LIV effects of $\tau=+400$ s TeV-1 and $\tau=-400$ s TeV-1, and offers a better view of the resolved pattern at high fluxes. The hysteresis pattern for X-ray energies (blue) is unaffected by LIV, as expected , and can be used as reference for a flat intrinsic trend at high flux. Conversely, the gamma-ray hysteresis pattern can be strongly modified (here in particular for the superluminal case with $\tau<0$) and in both cases shows at high fluxes a clear change in trend of the total time delays due to LIV effects, with the following behaviour:
  • subluminal LIV effects ($\tau>0$) induce an anticlockwise orientation of the hysteresis loop, and increasing time delays;
@@
Figure 9: HIDs in the gamma-ray energy domain for the flat trend case of Fig. 8 (centre), with various additional LIV contributions (in colour). In the inset the curves are shifted, emphasising the change in curvature orientation at high flux values.
Figure 10: HID highlighting hysteresis patterns for X-ray (blue) and gamma-ray (magenta) domains for the flat trend case shown in Fig. 8 (centre) with additional LIV effects, namely $\tau=+400$ s TeV-1 (top pannel) and $\tau=-400$ s TeV-1 (bottom pannel), showing that LIV effects can significantly modify the hysteresis pattern and even cause a change in loop orientation in the gamma-ray domain. While the X-ray hysteresis pattern is left unchanged relative to the case at $\tau=0$ in Fig. 8, a rotation pattern appears in the gamma-ray pattern (either clockwise or anticlockwise depending on the sign of $\tau$).
@@ 4 Conclusion and perspectives
This article was focused on energy-dependent time delays, arising as one of the most explored signatures of Lorentz invariance violation in astrophysics in the electromagnetic sector. Intrinsic time delays generated in situ by radiative processes of cosmic sources can interfere with these studies and need to be properly disentangled from propagation effects. It would indeed be necessary to accurately discriminate and quantify each contribution to any detected time delay to provide a legitimate constraint on the energy scale $E_{QG}$, and thus a proper interpretation on the tested quantum gravity models.
@@
We investigated this issue with a time dependent modelling of blazar flares induced by particle acceleration and relying on standard leptonic one-zone scenarios. Such models naturally generate intrinsic energy-dependent time delays whether considering SSC models, with or without adiabatic expansion or compression effects, or additional external inverse-Compton emission. The core of the present work focused on a multi-wavelength study that highlights a strong correlation and symmetry between intrinsic time delays in the X-ray (unaffected by LIV) and gamma-ray (where LIV effects could arise) energy domains. This correlation is due to the fact that in leptonic scenarios the same lepton population is at the origin of the two radiative processes. Two complementary approaches were then proposed to explore and quantify that correlation especially for the pure SSC case where it is the strongest, based on two observables, namely $(i)$ the Euclidean distance between the energy-dependent time delays at X-ray and very high energies, and $(ii)$ the hardness-intensity diagrams built from spectral energy distributions that show hysteresis patterns in the two domains. From these two methods, the presence of any independent external phenomenon modifying time delays such as LIV should be revealed by the weakening or even breaking of the expected intrinsic X-gamma correlation. Moreover, LIV should induce a systematic effect always affecting the correlation in the same direction, with an amplitude varying as a function of the redshift of the cosmic sources.
@@
The present study illustrates how all basic kinds of one-zone leptonic flares (SSC, with or without EIC and adiabatic effects) induced by particle acceleration can grow in two main regimes at high energies, either an acceleration-driven regime or a cooling-driven regime, depending on the high-energy lepton population evolution. The evolution of intrinsic time delays as a function of the energy follow three specific trends, described here as increasing, flat, and decreasing, which can be explained from the competition between acceleration and radiative processes. The flat trend corresponds to the thin transition between the two main regimes and is obtained only with a fine-tuning of the model parameters. EIC and adiabatic processes do not modify this general picture. However, their presence influences the nature of the flare compared to a pure SSC scenario, and can even change its regime. Additional EIC radiation increases radiative losses and always pushes the flares towards the decreasing trend and cooling-driven regime at VHE. Conversely, adiabatic expansion slows down particle acceleration processes and pushes the flares towards the increasing trend and the acceleration-driven regime (and the opposite for adiabatic compression). For these reasons, and for sake of simplicity, we mostly applied here the Euclidean distance and HID techniques to pure SSC flares. LIV effects acting only at VHE can reduce and even delete in some cases the strong correlation expected between the X-ray and VHE gamma-ray emission in leptonic scenarios. For a large domain of flare parameters, a minimum normalised Euclidean distance between delays in X-rays and in gamma-rays ($d_{E,\mathrm{min}}<0.6$) was deduced for SSC models of sources at low redshift (z = 0.03). Therefore, obtaining values of $d_{E,\mathrm{min}}>0.6$ from simultaneous X-ray and VHE observations of SSC flares from such sources would sign the presence of additional delays, possibly due to LIV effects. Similarly, the injection of LIV effects can significantly modify the HID hysteresis patterns at VHE energies, and can even invert the direction of rotation of the gamma-ray hysteresis curve compared to the X-ray curve. The simple detection of opposite directions of rotation of the hysteresis curve at high fluxes in X-rays and gamma-rays in SSC flares would indicate the presence of non-intrinsic time delays, possibly due to LIV.
@@
Euclidean distance and HID approaches may be more or less accurate and conclusive to discriminate between the origins of the delays, depending on the observed events and on the instrumental performance of the present or future detectors in terms of sensitivity, and spectral and temporal resolution. They both emphasise the decisive importance of simultaneous X-ray and VHE gamma-ray monitoring of variable blazars and of efficient global multi-wavelength alarm networks to catch both rise and decay phases of strong flares. Only under this condition will we succeed in disentangling possible LIV time delays from intrinsic ones with individual sources.
@@
The two techniques proposed here are closely related but complementary and can be adapted depending on the characteristics of available datasets. In particular, a high activity episode sometimes shows multiple flux peaks, such as in the case of the flare of blazar PKS 2155-304 in 2006 (Aharonian et al., 2007). Under these conditions, it is reasonable to assume that the multiple peaks will have similar intrinsic properties. The $d_{E}$ technique is based on the analysis of X-ray and gamma-ray light curves in several spectral bands. It should be possible to apply it to a single peak or to multiple peaks, as long as it is possible to measure the lag as a function of energy in the X-ray and VHE domains with a high enough accuracy. On the other hand, the HID method directly reflects the evolution of the SEDs and relies on the accurate measurement of the flux and spectral index variation in time. This accuracy will probably be achieved only by analysing individual peaks separately, provided the photon statistics are high enough. A first feasibility study for the use of HIDs with CTA is currently ongoing (Rosales De Leon et al., 2024).